# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 3.7 Position-Indexed Edge Formulation
As described in the previous section, the main advantage of EEF-CYCLE over the models described in Sections 3.3-3.5 is its polynomial model size. However, its LP-relaxation was theoretically shown to be relatively weak compared to CF-CYCLE and HCF-CYCLE (see Constantino et al., 2013 and Delorme et al., 2023). On the other hand, the cycle model PIEF-CYCLE, introduced by Dickerson et al., (2016), has both a polynomial model size and a tight LP-relaxation.
@@
We introduce a binary decision variable $x_{uv}^{sk}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in each position $k\in\mathcal{K}^{s}(u,v)$ of every subgraph $\mathcal{G}^{s}$, taking value $1$ if arc $(u,v)$ is selected at position $k$ of a cycle in subgraph $\mathcal{G}^{s}$, and 00 otherwise. This gives rise to the following model:
[…]
@@
The objective function (36) maximises the total weight and constraints (37) enforce that all RDPs are involved in at most one cycle. Constraints (38) are flow conservation constraints stating that for every graph copy, vertex, and position, it is only possible that a selected arc on that position leaves the vertex if an arc entering the vertex is selected on the previous position. In addition, by the definition of the sets $\mathcal{K}^{s}(u,v)$, these constraints enforce that if any arcs are selected from some subgraph $\mathcal{G}^{s}$, then these arcs form a cycle that involves vertex $s$.
Example 1.
(continued) Consider the instance presented in Figure 3 for $K=4$. The subgraphs required by PIEF-CYCLE are depicted in Figure 5, where the labels on the arcs are the sets $\mathcal{K}^{s}(u,v)$ of possible positions of the arcs in the subgraphs. In total there are $9$ variables, one for each feasible position of each arc in each subgraph. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}^{21}=x_{34}^{22}=x_{42}^{23}=1$.
$\{1\}$$\{2\}$$\{3\}$$\{2,4\}$$\{1\}$$\emptyset$
$\{1\}$$\{2\}$$\{3\}$
1234$\emptyset$
1234
@@
For each arc $(u,v)\in\mathcal{A}$, we consider a set $\mathcal{K}^{\prime}(u,v)$ of positions at which that arc can be selected in a chain. For all arcs $(u,v)\in\mathcal{A}_{\mathcal{N}}$ we have $\mathcal{K}^{\prime}(u,v)=\{1\}$, for all arcs $(u,v)\in\mathcal{A}_{\mathcal{R}}$ we have $\mathcal{K}^{\prime}(u,v)\subseteq\{2,\ldots,L-1\}$, and for all arcs $(v,\tau)\in\mathcal{A}_{\tau}$ we have $\mathcal{K}^{\prime}(v,\tau)=\{1\}$ if $v\in\mathcal{N}$ and $\mathcal{K}^{\prime}(v,\tau)\subseteq\{2,\ldots,L\}$, otherwise. We define a binary decision variable $y_{uv}^{k}$ for every arc $(u,v)\in\mathcal{A}$ and every possible position $k\in\mathcal{K}^{\prime}(u,v)$, taking value $1$ if arc $(u,v)$ is selected at position $k$ in a chain, and value 00 otherwise. Then, PIEF-CHAIN is defined as follows:
[…]
@@
The objective function (40) maximises the total weight and constraints (41) and (42) enforce that all RDPs and NDDs, respectively, are involved in at most one chain. Moreover, flow conservation constraints (43) enforce that for every vertex and every position, a selected arc may only leave the vertex in that position if an arc entering that vertex is selected in the previous position.
Example 1.
(continued) Consider the instance presented in Figure 3 for $L=4$. The graph required by PIEF-CHAIN is depicted in Figure 6, where the labels on the arcs are the sets $\mathcal{K}^{\prime}(u,v)$ of possible positions of the arcs. In total there are $20$ variables, one for each feasible position of each arc. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{51}^{1}=y_{12}^{2}=y_{23}^{3}=y_{3\tau}^{4}=1$.
[…]
Finally, we discuss improvements and techniques proposed for PIEF-CYCLE and PIEF-CHAIN. First, we comment on the construction of the sets $\mathcal{K}^{s}(u,v)$ in PIEF-CYCLE and $\mathcal{K}^{\prime}(u,v)$ in PIEF-CHAIN. Ideally, these sets are as small as possible to reduce the overall model size. However, computing the smallest possible sets is computationally demanding. Therefore, Dickerson et al., (2016) proposed an efficient method for constructing these sets using shortest path lengths, which we review in Appendix A.5. In addition, we present there a novel approach based on a breadth-first search algorithm, which results in smaller sets $\mathcal{K}^{s}(u,v)$ and $\mathcal{K}^{\prime}(u,v)$, and thus in a smaller model size, while having the same time complexity.
Moreover, Dickerson et al., (2016) showed that for PIEF-CYCLE, the variables for positions $1$ and $K$ can be eliminated when $K\geq 3$ and no self-loops are considered. Indeed, if the second arc in graph copy $\mathcal{G}^{s}$ leaves some vertex $u$, then arc $(s,u)$ must be chosen on position $1$, and if the $(K-1)$th arc in graph copy $\mathcal{G}^{s}$ goes to some vertex $v\neq s$, then arc $(v,s)$ must be selected on position $K$. For the sake of conciseness, we refer to Dickerson et al., (2016) for further details on this. Our preliminary experiments showed that omitting these variables only had a very minor impact on the empirical performance of this model, which is why we did not include this reduction in our implementation. Similarly, for PIEF-CHAIN, we can omit the variables associated to arcs $(v,\tau)\in\mathcal{A}_{\tau}$, which we describe in detail in Appendix A.6.
Furthermore, as for HCF-CYCLE and EEF-CYCLE, the vertex ordering plays a role for PIEFCYCLE. Dickerson et al., (2016) proposed to sort the vertices by descending total degree, as this resulted in a relatively small model size. The same rule was applied by Delorme et al., (2023), whereas Arslan et al., (2024) used a dynamic ordering rule.
@@ 4 Computational Experiments
One of our main goals was to empirically evaluate all of the model combinations resulting from combining the cycle and chain models discussed in Section 3. In Section 4.1 we outline the design of our experiments, in Section 4.2 we present the results of the experiments, and in Section 4.3 we provide a summary of those results.
@@ 4.1.1 Our Implementations of the Combined Models
We implemented all $5\times 7=35$ combinations of the cycle and chain models discussed in the previous sections, as well as the model EF-HYBRID. We summarise some properties of these models in Appendix A.4.
Whenever applicable, we applied graph reduction (as described in Appendix A.5) and we implicitly dealt with the terminal vertex $\tau$ (as described in Appendix A.6). Moreover, following the discussions throughout Section 3 regarding the importance of the ordering of the vertices for some of the cycle models, we re-indexed the vertices in $\mathcal{R}$ in descending order of degree for HCF-CYCLE and PIEF-CYCLE (to minimise its model size) and ascending order of degree for EEF-CYCLE (to benefit its LP relaxation). Furthermore, whenever a model includes an exponential number of constraints, we use constraint generation to handle the problematic constraints (as introduced in Section 3.5; see also Appendix A.4). In addition, preliminary experiments showed that the performance of our models depended on the method used by the ILP solver to solve the root node relaxation. We determined that the following rule works well: apply the primal simplex method for combined models involving EF-CYCLE (unless combined with HCF-CHAIN or EEF-CHAIN-EXP), EF-CHAIN-EXP or EEF-CHAIN-MTZ (unless combined with EEF-CYCLE), and apply the barrier method otherwise.
Furthermore, for the $3\times 2=6$ combined models obtained by combining cycle model CF-CYCLE, HCF-CYCLE or PIEF-CYCLE with chain model CF-CHAIN or PIEF-CHAIN, we also implemented a version that is solved through RCVF, following the framework by Delorme et al., (2023) and Delorme et al., (2024) (see also Section 3.3). We selected these models because they are characterised by a strong LP relaxation, which makes them particularly well-suited for RCVF. Moreover, as will be mentioned in Section 4.2.1, HCF-CYCLE and PIEF-CYCLE are the most effective among the five considered cycle models, while PIEF-CHAIN is the most effective chain model. Furthermore, CF-CYCLE and CF-CHAIN are included for their practical relevance, as these models allow for the modelling of certain objectives and constraints that cannot be handled using any of the other models (see e.g., Delorme et al., 2024). Our code is publicly available from the following open-source repository: https://doi.org/10.5281/zenodo.14905243 under a GNU GPL 3.0 licence.
@@ 4.1.2 Third-Party Methods
We tried to find and evaluate as many third-party methods as possible, focussing on methods based on advanced techniques such as B&P. Our coverage includes implementations that are publicly available, or for which we were able to obtain the code directly from the authors, but some methods were introduced too recently to be included in the evaluation. Also, the constraint programming approaches referred to in Section 2.1 were not considered as they do not tackle KE-Opt directly. The third-party methods that we did include in our experiments are summarised in Table 2, where they are classified according to their underlying combined model. All of these methods, except CG-TSP, use B&P and were discussed in Section 3.3. CG-TSP is a Java implementation of the combined model EF-CYCLE+EEF-CHAIN-EXP, which like our C++ implementation of that model, uses constraint generation. For each third-party method, we indicate in the table whether the code is available in a public repository (linked either in the accompanying paper or on the author’s website), or whether the code was obtained directly from the authors. Furthermore, for each third-party method we indicate its programming language and ILP solver.
@@
EF-CYCLE CG-TSP Anderson et al., (2015) Public Java CPLEX 22.1.1
@@
$\dagger$) The version of BNP-DFS that is tested here is associated with Abraham et al., (2007),
 but it is not the most recent version of their code.
[…]
 Furthermore, it does not apply to weighted instances and neither does it allow for chains of length 1.
*) JL-BNP and JL-BNP-PICEF support multiple solvers; we list here the solver used in our experiments.
@@ 4.1.3 Instance Generation
We generated compatibility graphs using the instance generator created by Delorme et al., (2022), which is available at https://wpettersson.github.io/kidney-webapp/#/generator. We used their “SplitPRA BandXMatch PRA0” profile, which was shown by Delorme et al., (2022), to create instances with similar characteristics to those found in historical datasets from the UK’s national KEP. These compatibility graphs do not contain self-loops, and each recipient has a single donor. For each ||$|\mathcal{R}|\in\{50,100,200,500,750,1000\}$ and $|\mathcal{N}|\in\{0.05|\mathcal{R}|,0.10|\mathcal{R}|,0.20|\mathcal{R}|\}$, we created ten random graphs with $|\mathcal{R}|$ RDPs and $|\mathcal{N}|$ NDDs. Moreover, for each graph we considered a weighted and an unweighted variant. In both cases, the weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ were set to 0. The remaining weights $w_{uv}$ for arcs $(u,v)\in\mathcal{A}_{\mathcal{N}}\cup\mathcal{A}_{\mathcal{R}}$ were randomly sampled from the discrete uniform distribution on the set $\{1,2,\ldots,91\}$ in the weighted case [^7] The generator that we used samples weights from $\{0,1,\ldots,90\}$, but we add $1$ to the weight of each arc to avoid having arcs with a weight of 00, reflecting that all transplants have a positive utility., and set to $1$ in the unweighted case. In addition, we considered cycle length limits $K\in\{3,4,5,6\}$ and chain length limits $L\in\{K,K+1,2K\}$. This resulted in a total of $6\times 3\times 10\times 2\times 4\times 3=4320$ instances. These instances are publicly available from the following data repository: https://doi.org/10.5525/gla.researchdata.1878, under a CC BY 4.0 licence.
@@ 4.2 Computational Results
The full results of our simulations can be found on the following webpage: https://www.optimalmatching.com/kep-survey-2025/, where one can easily create custom heatmaps for varying subsets of instances, methods and performance indicators. In this paper, we focus on some of the more interesting results, which we present in the following four parts. First, in Section 4.2.1 we present the overall results across all instances of our standard implementations of the combined models in order to evaluate which models are most effective in terms of the number of optimal solutions found and average running time, providing justification where appropriate. Second, in Section 4.2.2 we do the same, but for the RCVF implementations of the models where we implemented this. Third, in Section 4.2.3 we focus on the overall performance of the third-party methods under the same set-up. Fourth, in Appendix A.8 we evaluate the most effective methods (from both our implementations and the third-party methods) on different subsets of the instances to understand how different instance parameters influence the behaviour of the methods as well as the computational hardness of KE-Opt.

### v2

@@ 3.7 Position-Indexed Edge Formulation
As described in the previous section, the main advantage of EEF-CYCLE over the models described in Sections 3.3-3.5 is its polynomial model size. However, its LP relaxation was theoretically shown to be relatively weak compared to CF-CYCLE and HCF-CYCLE ( see Section 3.8). On the other hand, the cycle model PIEF-CYCLE, introduced by Dickerson et al., (2016), has both a polynomial model size and an LP relaxation that is as strong as that of CF-CYCLE and HCF-CYCLE ( see Section 3.8).
@@
We introduce a binary decision variable $x_{uv}^{sk}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in each position $k\in\mathcal{K}^{s}(u,v)$ of every subgraph $\mathcal{G}^{s}$, taking value $1$ if arc $(u,v)$ is selected at position $k$ of a cycle in subgraph $\mathcal{G}^{s}$, and 0 otherwise. This gives rise to the following model:
$(\texttt{PIEF-CYCLE})\quad$ $\max$ $\sum_{s\in\mathcal{R}}\sum_{(u,v)\in\mathcal{A}^{s}}\sum_{k\in\mathcal{K}^{s}(u,v)}w_{uv}x_{uv}^{sk}$ (36)
@@
The objective function (36) maximises the total weight and constraints (37) enforce that each RDP is involved in at most one cycle. Constraints (38) are flow conservation constraints stating that for every graph copy, vertex, and position, it is only possible that a selected arc on that position leaves the vertex if an arc entering the vertex is selected on the previous position. In addition, by the definition of the sets $\mathcal{K}^{s}(u,v)$, these constraints enforce that if any arcs are selected from some subgraph $\mathcal{G}^{s}$, then these arcs form a cycle that involves vertex $s$.
@@
For each arc $(u,v)\in\mathcal{A}$, we consider a set $\mathcal{K}^{\prime}(u,v)$ of positions at which that arc can be selected in a chain. For each arc $(u,v)\in\mathcal{A}_{\mathcal{N}}$ we have $\mathcal{K}^{\prime}(u,v)=\{1\}$, for each arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$ we have $\mathcal{K}^{\prime}(u,v)\subseteq\{2,\ldots,L-1\}$, and for each arc $(v,\tau)\in\mathcal{A}_{\tau}$ we have $\mathcal{K}^{\prime}(v,\tau)=\{1\}$ if $v\in\mathcal{N}$ and $\mathcal{K}^{\prime}(v,\tau)\subseteq\{2,\ldots,L\}$, otherwise. We define a binary decision variable $y_{uv}^{k}$ for every arc $(u,v)\in\mathcal{A}$ and every possible position $k\in\mathcal{K}^{\prime}(u,v)$, taking value $1$ if arc $(u,v)$ is selected at position $k$ in a chain, and value 0 otherwise. Then, PIEF-CHAIN is defined as follows:
$(\texttt{PIEF-CHAIN})\quad$ $\max$ $\sum_{(u,v)\in\mathcal{A}}\sum_{k\in\mathcal{K}^{\prime}(u,v)}w_{uv}y_{uv}^{k}$ (40)
@@
The objective function (40) maximises the total weight and constraints (41) and (42) enforce , respectively, that each NDD and each RDP is involved in at most one chain. Moreover, flow conservation constraints (43) enforce that for every vertex and every position, a selected arc may only leave the vertex in that position if an arc entering that vertex is selected in the previous position.
Finally, we discuss improvements and techniques proposed for PIEF-CYCLE and PIEF-CHAIN. First, we comment on the construction of the sets $\mathcal{K}^{s}(u,v)$ in PIEF-CYCLE and $\mathcal{K}^{\prime}(u,v)$ in PIEF-CHAIN. Ideally, these sets are as small as possible to reduce the overall model size. However, computing the smallest possible sets is computationally demanding. Therefore, Dickerson et al., (2016) proposed an efficient method for constructing these sets using shortest path lengths, which we review in Appendix F. In addition, we present there a novel approach based on a breadth-first search algorithm, which results in smaller sets $\mathcal{K}^{s}(u,v)$ and $\mathcal{K}^{\prime}(u,v)$, and thus in a smaller model size, while having the same time complexity.
Moreover, Dickerson et al., (2016) showed that for PIEF-CYCLE, the variables for positions $1$ and $K$ can be eliminated when $K\geq 3$ and no self-loops are considered. Indeed, if the second arc in graph copy $\mathcal{G}^{s}$ leaves some vertex $u$, then arc $(s,u)$ must be chosen on position $1$, and if the (K1)(K-1)th arc in graph copy $\mathcal{G}^{s}$ goes to some vertex $v\neq s$, then arc $(v,s)$ must be selected on position $K$. For the sake of conciseness, we refer to Dickerson et al., (2016) for further details on this. Our preliminary experiments showed that omitting these variables only had a very minor impact on the empirical performance of this model, which is why we did not include this reduction in our implementation. Similarly, for PIEF-CHAIN, we can omit the variables associated to arcs $(v,\tau)\in\mathcal{A}_{\tau}$, which we describe in detail in Appendix G.
Furthermore, as for HCF-CYCLE and EEF-CYCLE, the vertex ordering plays a role for PIEF-CYCLE. Dickerson et al., (2016) proposed to sort the vertices by descending total degree, as this resulted in a relatively small model size. The same rule was applied by Delorme et al., (2023), whereas Arslan et al., (2024) used a dynamic ordering rule.
@@
3.8 Comparison of Theoretical Properties
There are several known results about the LP relaxations of the cycle and chain models. First, for the cycles-only case of KE-Opt, Constantino et al., (2013) showed that CF-CYCLE dominates both EF-CYCLE and EEF-CYCLE in terms of LP relaxation, while there is no dominance relation between EF-CYCLE and EEF-CYCLE. Subsequently, Dickerson et al., (2016) and Delorme et al., (2023) proved that the LP relaxations of PIEF-CYCLE and HCF-CYCLE, respectively, are as strong as that of CF-CYCLE. When chains are also considered, the only known result is one by Dickerson et al., (2016), stating that the LP relaxation of CF-CYCLE+CF-CHAIN dominates that of CF-CYCLE+PIEF-CHAIN. We summarise some further properties of the considered models in Appendix E.
@@ 4 Computational Experiments
One of our main goals was to empirically evaluate all of the model combinations resulting from combining the cycle and chain models discussed in Section 3. In Section 4.1 we outline the design of our experiments, in Section 4.2 we present the results of the experiments, and in Section 4.3 we summarise those results.
@@ 4.1.1 Our Implementations of the Combined Models
We implemented all $5\times 7=35$ combinations of the cycle and chain models discussed in the previous sections, as well as the model EF-HYBRID. Whenever applicable, we applied graph reduction (as described in Appendix F) and we implicitly dealt with the terminal vertex $\tau$ (as described in Appendix G). Moreover, following the discussions throughout Section 3 regarding the importance of the ordering of the vertices for some of the cycle models, we re-indexed the vertices in $\mathcal{R}$ in descending order of degree for HCF-CYCLE and PIEF-CYCLE (to minimise its model size) and ascending order of degree for EEF-CYCLE (to benefit its LP relaxation). Furthermore, whenever a model includes an exponential number of constraints, we use constraint generation to handle the problematic constraints (as introduced in Section 3.5; see also Appendix E). In more detail, in our implementation of constraint generation, we only check potential incumbents (i.e., feasible integer solutions) for violated constraints, and the added constraints apply globally across the entire branch-and-bound tree. In addition, preliminary experiments showed that the performance of our models depended on the method used by the ILP solver to solve the root node relaxation. We determined that the following rule works well: apply the primal simplex method for combined models involving EF-CYCLE (unless combined with HCF-CHAIN or EEF-CHAIN-EXP), EF-CHAIN-EXP or EEF-CHAIN-MTZ (unless combined with EEF-CYCLE), and apply the barrier method otherwise.
Furthermore, for the $3\times 2=6$ combined models obtained by combining cycle model CF-CYCLE, HCF-CYCLE or PIEF-CYCLE with chain model CF-CHAIN or PIEF-CHAIN, we also implemented a version that is solved through RCVF, following the framework by Delorme et al., (2023) and Delorme et al., (2024) (see also Section 3.3). We selected these models because they have the strongest LP relaxation, which makes them particularly well-suited for RCVF. Moreover, as will be mentioned in Section 4.2.1, HCF-CYCLE and PIEF-CYCLE are the most effective among the five considered cycle models, while PIEF-CHAIN is the most effective chain model. Furthermore, CF-CYCLE and CF-CHAIN are included for their practical relevance, as these models allow for the modelling of certain objectives and constraints that cannot be handled using any of the other models (see e.g., Delorme et al., 2024). Our code is publicly available from the following open-source repository: https://doi.org/10.5281/zenodo.14905243 under a GNU GPL 3.0 licence.
@@ 4.1.2 Third-Party Methods
We tried to find and evaluate as many third-party methods as possible, focussing on methods based on advanced techniques such as B&P. Our coverage includes implementations that are publicly available, or for which we were able to obtain the code directly from the authors, but some methods were introduced too recently to be included in the evaluation. Also, the constraint programming approaches referred to in Section 2.1 were not considered as they do not tackle KE-Opt directly. The third-party methods that we did include in our experiments are summarised in Table 2, where they are classified according to their underlying combined model. All of these methods, except CG-TSP, use B&P and were discussed in Section 3.3. CG-TSP is a Java implementation of the combined model EF-CYCLE+EEF-CHAIN-EXP, which like our C++ implementation of that model, uses constraint generation. However, their algorithm also adds constraints that are violated by fractional solutions. For each third-party method, we indicate in the table whether the code is available in a public repository (linked either in the accompanying paper or on the author’s website), or whether the code was obtained directly from the authors. Furthermore, for each third-party method we indicate its programming language and ILP solver.
@@
EF-CYCLE CG-TSP Anderson et al., 2015b Public Java CPLEX 22.1.1
@@
$\dagger$) The version of BNP-DFS that is tested here is associated with Abraham et al., (2007), but it is not the most recent version of their code. $\ddagger$}) DCD is based on the chain-to-cycle transformation (see Section 3.2), meaning that it is limited to scenarios where $K=L$. Furthermore, it does not apply to weighted instances and neither does it allow for chains of length 1. *) JL-BNP and JL-BNP-PICEF support multiple solvers; we list here the solver used in our experiments.
@@ 4.1.3 Instance Generation
In our experiments we focus on instances that reflect well the characteristics of real KEP datasets. Therefore, we generated compatibility graphs using one of the state-of-the-art static generators mentioned in Section 2.8, namely the instance generator created by Delorme et al., (2022), which is available at https://wpettersson.github.io/kidney-webapp/#/generator. We used their “SplitPRA BandXMatch PRA0” profile, which was shown by Delorme et al., (2022) to create instances with similar characteristics to those found in historical datasets from the UK’s national KEP. These compatibility graphs do not contain self-loops, and each recipient has a single donor. For each $|\mathcal{R}|\in\{50,100,200,500,750,1000\}$ and $|\mathcal{N}|\in\{0.05|\mathcal{R}|,0.10|\mathcal{R}|,$ $0.20|\mathcal{R}|\}$, we created ten random graphs with $|\mathcal{R}|$ RDPs and $|\mathcal{N}|$ NDDs. Moreover, for each graph we considered a weighted and an unweighted variant. In both cases, the weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ were set to 0 to ensure compatibility with some of the third-party methods. The remaining weights $w_{uv}$ for arcs $(u,v)\in\mathcal{A}_{\mathcal{N}}\cup\mathcal{A}_{\mathcal{R}}$ were randomly sampled from the discrete uniform distribution on the set $\{1,2,\ldots,91\}$121212 The generator of Delorme et al., (2022) samples weights from $\{0,1,\ldots,90\}$. We modify this by adding 1 to each weight generated to avoid transplants with non-positive weight. in the weighted case, and set to $1$ in the unweighted case. In addition, we considered cycle length limits $K\in\{3,4,5,6\}$ and chain length limits $L\in\{K,K+1,2K\}$. This resulted in a total of $6\times 3\times 10\times 2\times 4\times 3=4320$ instances. These instances are publicly available from the following data repository: https://doi.org/10.5525/gla.researchdata.1878, under a CC BY 4.0 licence. To put our instances into perspective, in recent years the UK’s KEP has had about 250250-$300$ RDPs and up to $30$ NDDs per matching run. While this programme is one of the largest in the world, it allows only cycles and chains of length up to 3. In contrast, some other countries support larger exchanges: for example, the Dutch KEP permits cycles and chains of length 4, while the Czech and Spanish programs have allowed chains of length up to 6 (Biró et al., , 2021).
@@ 4.2 Computational Results
The full results of our simulations can be found on the following webpage: https://www.optimalmatching.com/kep-survey-2025/, where one can easily create custom heatmaps for varying subsets of instances, methods and performance indicators. In this paper, we focus on some of the more interesting results, which we present in the following four parts. First, in Section 4.2.1 we present the overall results across all instances of our standard implementations of the combined models in order to evaluate which models are most effective in terms of the number of optimal solutions found and average running time, providing justification where appropriate. Second, in Section 4.2.2 we focus on the overall performance of the third-party methods under the same set-up. Third, in Appendix I we do the same, but for the RCVF implementations of the models where we implemented this. Fourth, in Appendix J we evaluate the most effective methods (from both our implementations and the third-party methods) on different subsets of the instances to understand how different instance parameters influence the behaviour of the methods as well as the computational hardness of KE-Opt.